OpenStudy (anonymous):

suppose f(x)=x^3-3 find the graph of f(1/3x)

OpenStudy (misty1212):

HI!!\[f(\frac{1}{3x})\]?

OpenStudy (misty1212):

of \[f(\frac{1}{3}x)\]?

OpenStudy (anonymous):

the x is to the side of the fraction

OpenStudy (misty1212):

ok same as \[f(\frac{x}{3})\] then

OpenStudy (misty1212):

graph \[y=\left(\frac{x}{3}\right)^3-3\]

OpenStudy (misty1212):

or \[y=\frac{x^3}{27}-3\] same thing